OpenStudy (anonymous):

Gary has a length of string for making a yo-yo that is fifty-seven and one-half inches long. He wants to cut that in half to get the proper length. Which expression can be used to find the length of each piece of string? one-half times fifty-seven and one-half two times fifty-seven and one-half fifty-seven and one-half divided by one-half two divided by fifty-seven and one-half

OpenStudy (anonymous):

@HatcrewS @happycomposers @Prototype93 @tkhunny @berrypicking

OpenStudy (tkhunny):

Still tagging the world, eh? Please stop it. Just post your problem and SHOW YOUR WORK! What do you think? Where did you land? Why do you believe you did not manage a correct response?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

OK. first of all I posted my question like 5 minutes ago no one responded. IF you do not want to answer then go else where this IS a site people get help with their homework on So you don't have to be a jerk about it. By the way I do show my OWN work. K thx.

OpenStudy (anonymous):

http://gyazo.com/97d77215d0cc2126fa478b5932c52845 It's a multiple choice question that i do NOT understand how to do??.. How should i "SHOW MY WORK"?

OpenStudy (tkhunny):

Okay, I'll give you one more free one. Here's how I might show my work. Truthfully, I do not recall having been presented a problem quite like this one. one-half times fifty-seven and one-half two times fifty-seven and one-half fifty-seven and one-half divided by one-half two divided by fifty-seven and one-half Okay, those are my choices. Let's see... The string is 57½. I need to divide that into two pieces. Okay, that means three things: 1) I will need a division problem. This discards the first 2. 2) I am dividing into two parts, so I will need a 2 in the denominator. Whoops! We don't have one of those. The correct answer isn't there! That's alarming! 3) The original string is 57½, so that will have to be in the numerator. This discards the last one. Well, the third one is the closest, but it's still wrong. We need 57½/2, not 57½/(1/2)! Wait just a minute! Multiplication and division are kind of the same thing. Myltiplying by 1/2 is the same as dividing by 2. I better rewrite those expressions and make sure. one-half times fifty-seven and one-half -- That's \(\dfrac{1}{2}\cdot 57½ = \dfrac{57½}{2}\) -- This one actually looks pretty good, now. two times fifty-seven and one-half -- That's \(2\cdot 57½\) -- Still no good. fifty-seven and one-half divided by one-half -- That's \(\dfrac{57½}{\dfrac{1}{2}}\) -- No good. That denominator is just wrong. two divided by fifty-seven and one-half -- That's \(\dfrac{2}{57½}\) -- No good. That's upside down. There we have it. It was a tricky little fellow, but once we figured out that division problems can come from multiplication problems, we figured it out. The first one is the right way to go about the division of the string. It is a strangely unknown fact that problem sets in math books relate directly to the material just presented. You are very, VERY unlikely to be presented with a problem that you cannot solve with just a little time and energy thinking about the problem rather than quickly giving up. The fact that you said, "I do NOT understand how to do." suggests to me that you expect some magical understanding fairy to drop from the sky. That doesn't happen. The idea is to cause you to think in what may be a new way. Waiting for understanding isn't likely to be beneficial. Thinking in an orderly fashion is likely to be beneficial.

OpenStudy (anonymous):

Um? you could've just told me that in the first place. News flash .. I'm already finished with my test. You responded way TOO late Sorry dear. </3

OpenStudy (tkhunny):

It is only too late if you fail to learn. Or maybe too early. When you want to learn, you will find that listening to me (or any other capable instructor of mathematics) is a good idea. As you continue only to argue and resist, it is very difficult to help you. I could not have just told you that in the first place, since that would be in direct violation of the Code of Conduct. You really should read it and heed its wisdom. I do not understand why you resist so strenuously.

OpenStudy (anonymous):

._. well then..